OpenStudy (anonymous):

Find the partial fraction decomposition for f(x)= (28)/(x^2(x+2)(x^2-x+1))

OpenStudy (anonymous):

\[ -\frac{4 (2 x+1)}{x^2-x+1}+\frac{14}{x^2}+\frac{1}{x+2}+\frac{7}{x} \]